# Triple-BERT: Do We Really Need MARL for Order Dispatch on Ride-Sharing Platforms?

### v1

@@ Abstract
On-demand ride-sharing platforms, such as Uber and Lyft, face the intricate real-time challenge of bundling and matching passengers—each with distinct origins and destinations—to available vehicles, all while navigating significant system uncertainties. Due to the extensive observation space arising from the large number of drivers and orders, order dispatching, though fundamentally a centralized task, is often addressed using Multi-Agent Reinforcement Learning (MARL). However, independent MARL methods fail to capture global information and exhibit poor cooperation among workers, while Centralized Training Decentralized Execution (CTDE) MARL methods suffer from the curse of dimensionality. To overcome these challenges, we propose Triple-BERT, a centralized Single Agent Reinforcement Learning (MARL) method designed specifically for large-scale order dispatching on ride-sharing platforms. Built on a variant TD3, our approach addresses the vast action space through an action decomposition strategy that breaks down the joint action probability into individual driver action probabilities. To handle the extensive observation space, we introduce a novel BERT-based network, where parameter reuse mitigates parameter growth as the number of drivers and orders increases, and the attention mechanism effectively captures the complex relationships among the large pool of driver and orders. We validate our method using a real-world ride-hailing dataset from Manhattan. Triple-BERT achieves approximately an 11.95% improvement over current state-of-the-art methods, with a 4.26% increase in served orders and a 22.25% reduction in pickup times. Our code, trained model parameters, and processed data are publicly available at the repository https://github.com/RS2002/Triple-BERT.  [^1] Do We Really Need MARL for Order Dispatch on Ride-Sharing Platforms? In summary: No, because we developed a SARL method that achieves better global planning. However, yes, our method still requires MARL for pre-training to establish a strong starting point for SARL.
@@ 1 Introduction
Recently, Reinforcement Learning (RL) methods have shown great potential in addressing the order dispatching problem in ride-sharing platforms. Model-free RL, in particular, enables agents to autonomously learn optimal dispatching policies by interacting with the environment, without requiring complex system modeling. This approach allows platforms to optimize multiple objectives, including platform income, driver payments, and customer satisfaction. Despite these advantages, applying RL to large-scale order dispatching introduces significant challenges. The vast action and observation spaces, stemming from the large number of drivers and orders, make sufficient exploration and efficient training difficult. Multi-Agent Reinforcement Learning (MARL) methods have been widely adopted to address these challenges by decomposing the problem into smaller subproblems for individual agents (drivers). Independent MARL methods, such as Independent Double DQN (IDDQN) [1; 17; 31] and Independent SAC (ISAC) [37], are computationally efficient but fail to capture global information and exhibit limited cooperation among agents. Graph Neural Networks (GNNs) have been introduced to enable the network to capture neighboring information for each agent, alleviating this issue to certain extent [16; 34]. Meanwhile, Centralized Training with Decentralized Execution (CTDE) methods, such as QMIX [12] and Coordinated Policy Optimization (CoPO) [32], struggle with the curse of dimensionality when applied to large-scale scenarios with thousands of agents, resulting in slow convergence and suboptimal performance.
@@
To address these limitations, this paper proposes a centralized Single-Agent Reinforcement Learning (SARL) method, named Triple-BERT, tailored for large-scale order dispatching in ride-sharing platforms. Triple-BERT introduces an action decomposition method that simplifies the joint action probability into individual driver action probabilities, enabling each driver to make independent decisions while maintaining global coordination. The method leverages TD3 [10] for optimization, with modifications to the actor optimization process via policy gradient [26] to better suit the ride-sharing context. To handle the extensive observation space, we design a novel BERT-based [7] neural network architecture. This network employs bi-directional self-attention to effectively capture complex relationships between drivers and orders, while its parameter reuse mechanism prevents parameter explosion as the number of drivers and orders increases. Additionally, compared to MARL, SARL faces a unique challenge of sample scarcity, as the records of multiple agents are merged into a single training stream. To address this, we propose a two-stage training strategy, where feature extractors are pre-trained using a MARL approach to learn general embedding capabilities, followed by centralized fine-tuning. The main contributions of this paper can be summarized as follows:
@@
  • We introduce Triple-BERT, which is the first centralized SARL framework for large-scale order dispatching on ride-sharing platforms. This approach addresses the limitations of the observation space and the inefficiencies in cooperation among agents present in MARL methods. To tackle the large action space inherent in the matching problem of order dispatching tasks, we propose an action decomposition method that breaks down the joint action probability into individual driver action probabilities. Additionally, we propose a two-stage training method to address the sample scarcity issue in SARL, where the feature extractors are first trained using a MARL approach.
@@ 2 Problem Setup
Figure 1: Workflow
@@ 3 Methodology
Figure 2: Network Architecture
@@ 3.1 Overview
In this work, we aim to utilize centralized SARL to address the large-scale order dispatching task, with the goal of enabling the model to fully leverage global information to enhance cooperation among workers. To tackle the challenges of large action and observation spaces, we propose a novel network architecture, as illustrated in Fig. 2. This architecture employs the BERT model [7] to effectively extract the relationships between workers and orders using the self-attention mechanism. Additionally, an improved QK-attention [39] is implemented to reduce the computational complexity associated with the order dispatching task. Furthermore, we introduce an action decomposition method that breaks down the choice probability of each action within the vast action space into individual action probabilities for each worker selecting each order. Finally, to address the data scarcity challenge in MARL, we propose a two-stage training method, as shown in Fig. 1. In the first stage, we train the upstream layers of the network using the IDDQN approach, allowing them to develop general feature extraction capabilities. Subsequently, we train the entire neural network using centralized TD3 to realize better cooperation between workers.
@@ 3.2.1 Feature Extractors
At each time step, the network takes the entire state $S_{t}=[W_{t},O_{t}]$ as input. We consider this as a combination of two sequences: $W_{t}$ and $O_{t}$. For each element $w_{i,t}$ and $o_{j,t}$, we employ two distinct encoders, referred to as the ”Worker Encoder” and the ”Order Encoder”, to embed them separately into a feature space of the same dimension, allowing them to be input into a single BERT model.
Each worker state $w_{i,t}$ consists of two parts: an on-board order sequence and other non-sequence information. For the order sequence, a bi-directional LSTM [13] is utilized to extract its features. This approach effectively encodes variable-length sequences into a uniform dimensional feature space, addressing the curse of dimensionality associated with conventional MLP encoders, where the number of parameters increases with sequence length. For non-sequence information, an MLP is employed for feature extraction. Finally, the two features are combined into a primary feature $\tilde{w}_{i,t}$. For the orders to be assigned $o_{j,t}$, an MLP is also used to extract the feature $\tilde{o}_{j,t}$. Notably, the dimensions of $\tilde{w}_{i,t}$ and $\tilde{o}_{j,t}$ are identical, and their information is concatenated into a sequence represented as $\tilde{S}_{t}=[\tilde{w}_{1,t},\tilde{w}_{2,t},\ldots,\tilde{w}_{n,t},\tilde{o}_{1,t},\tilde{o}_{2,t},\ldots,\tilde{o}_{m_{t},t}]$.
Additionally, to facilitate network convergence and enhance the extraction of input features, we incorporate a normalization layer and an Adaptive Re-weighting Layer (ARL) [4]. Given that different parts of the input may have varying magnitudes, which can impede model training, the normalization layer effectively addresses this issue. Furthermore, since different parts of the input carry different levels of importance, we utilize the ARL to enable the model to learn these variations, represented as: $y=x\circ\Omega$, where $x$ denotes the input, $\Omega$ represents the weight vector, calculated by $\Omega=\text{MLP}(x)$, and $\circ$ indicates the element-wise product.
@@ 3.2.2 Actor Sub-Networks
The Actor sub-network consists of a BERT [7] model for feature extraction and a QK-attention module [39] for action decomposition and generation, which we will introduce in turn. In the feature extractors, we have already extracted the primary features from each worker and order state separately. To further explore the relationships between workers and orders, we utilize the BERT model, where the self-attention mechanism can effectively capture these relationships: $\overline{S}_{t}=[\overline{w}_{1,t},\overline{w}_{2,t},\ldots,\overline{w}_{n,t},\overline{o}_{1,t},\overline{o}_{2,t},\ldots,\overline{o}_{m_{t},t}]=\text{Actor-BERT}(\tilde{S}_{t})$. Specifically, due to the permutation invariance of our input sequence, we omit the positional embedding in BERT, ensuring that the order in $S$ does not influence the encoding result. In contrast to conventional MARL methods like [16; 12], which encode each worker with its neighboring states to gain a broader perspective, our Actor-BERT directly aggregates global worker information, facilitating more effective cooperative dispatching between workers.
In conventional order dispatching tasks, the typical approach to address the dynamic action space (related to the number of orders) involves evaluating each worker-order pair separately and finding the optimal dispatching solution based on these evaluations. However, this approach has two significant shortcomings. First, it neglects the relationships between orders, which we address through the self-attention mechanism in BERT, capturing not only the relationships between workers but also between orders and between orders and workers. Second, evaluating each worker-order pair is time-consuming and resource-intensive: $\text{F}(\overline{w}_{i,t},\overline{o}_{j,t};\theta_{F})\in\mathbb{R}^{1}$, where F is the network and $\theta_{F}$ represents its parameters. The complexity can be represented as $O(|\text{F}|\cdot n\cdot m_{t})$, where $|\text{F}|$ denotes the complexity of the neural network. To mitigate this issue, we employ a QK-attention module [39], represented as:
@@
where f and g are two smaller networks, and $\theta_{f}$ and $\theta_{g}$ are their parameters. The intuition behind QK-attention is to use two smaller networks to approximate a larger network, similar to the motivation behind LoRA [14]. In this way, the complexity of computing all worker-order pairs becomes $O(|\text{f}|\cdot n+|\text{g}|\cdot m_{t}+d\cdot n\cdot m_{t})$, where $|\text{f}|$ and $|\text{g}|$ are the complexities of the two neural networks, $d$ is their output dimension, and $d\cdot n\cdot m_{t}$ is the complexity of matrix multiplication. Here, $d$ is very small, making $d\cdot n\cdot m_{t}$ much smaller than the neural network computation complexity, i.e., $d\cdot n\cdot m_{t}\ll|\text{f}|\approx|\text{g}|<|\text{F}|$. Thus, we have $O(|\text{f}|\cdot n+|\text{g}|\cdot m_{t}+d\cdot n\cdot m_{t})<O(|\text{F}|\cdot(n+m_{t}))<O(|\text{F}|\cdot n\cdot m_{t})$, indicating that the QK-attention successfully transforms the multiplication complexity of evaluating each worker-order pair into addition complexity.
However, we observe a parameter redundancy issue in Equation 2, which can lead to potential instability during training. This redundancy arises because there are actually infinite solutions for $f$ and $g$, as $f^{\prime}=\alpha f$ and $g^{\prime}=\frac{g}{\alpha}$ is also a valid solution for any non-zero real vector $\alpha$. Inspired by Dueling DQN [35], we propose a positive normalization method:
@@ 3.2.3 Critic Sub-Networks
where $\overline{w}_{i,t}$ and $\overline{o}_{j,t}$ are the outputs of Actor-BERT, and $(\overline{w}_{i,t},\overline{o}_{j,t})$ represents the combination of the two vectors into a single feature vector. We then construct a new sequence: $\dot{S}_{t}=[\mathcal{A}(w_{1,t}),\mathcal{A}(w_{2,t}),\ldots,\mathcal{A}(w_{i,t})]$. Another BERT network, referred to as ”Critic-BERT”, is used to further extract features from $\dot{S}_{t}$, represented as $\ddot{S}_{t}=\text{Critic-BERT}(\dot{S}_{t})$. A self-attention mechanism and a linear layer (collectively named Critic-MLP) are then utilized to estimate the Q-value from $\ddot{S}_{t}$ (for detailed processing methods, refer to [6]). Furthermore, as TD3 [10] requires two critics, we employ two distinct Critic-BERT and Critic-MLP networks. These share the input features from Actor-BERT but process them separately.
@@ 3.3.1 Stage 1: Decentralized IDDQN Training
To construct the network, we utilize QK-attention to process the outputs of the worker encoder and order encoders to estimate the Q-value for each worker-order pair, represented as $\text{QK-Attention-Norm}(\tilde{w}_{i,t},\tilde{o}_{j,t})$ (denoted as $y_{i,j,t}$). Although the state space encompasses the entire order state from $o_{1,t}$ to $o_{m_{t},t}$, we focus on a single order $o_{j,t}$ when computing the Q-value for choosing order $j$. This approach aligns with previous work such as [16; 15], as the entire order state can be excessively large for a simple network to learn (our Triple-BERT effectively addresses this issue) and many networks struggle to process variable dimensional inputs (with order amounts varying at each time step). Consequently, we can compute a Q-matrix $Y_{t}\in\mathbb{R}^{n,m_{t}}$, where the element in the $i$-th row and $j$-th column, $y_{i,j,t}$, represents the Q-value of assigning order $j$ to worker $i$ at time $t$. The core strategy of IDDQN is to maximize the global Q-value, expressed as $\text{Q}(S_{t},A_{t})=\sum_{i=1}^{n}\text{Q}(s_{i,t},a_{i,t})$ at each time step. To achieve this, we construct a bipartite graph where each worker and order is represented as a node. An arbitrary worker $i$ and order $j$ are linked by an edge weighted by the Q-value of this worker selecting this order at the current time, i.e., $y_{i,j,t}$. We then utilize Integer Linear Programming (ILP) to solve this maximizing bipartite matching problem. (To avoid assigning orders to unavailable workers—those at full capacity or on their way to pick up an assigned order—we set the Q-value of all actions for such workers in the Q-matrix $Y_{t}$ to -$\infty$.) A detailed construction of the problem is provided in Appendix B.1. For the training of IDDQN, it follows the same process of previous work [16]. Due to page limitation, we detailed it in Appendix D.1.
Here, We want to make some explanations about the independent assumption. We acknowledge that the independent assumption can be unreliable, which indeed hinders the performance of conventional independent MARL-based methods. However, previous works [24; 9; 33] have shown the efficacy and simplicity of independent MARL methods. As a result, even if the performance of independent MARL is not satisfactory, it does provide a good starting point for our centralized SARL method. In our approach, the independent assumption is only utilized during the pre-training stage to warm up the model. After pre-training, our centralized SARL framework no longer relies on the independence assumption.

### v2

@@ Abstract
On-demand ride-sharing platforms, such as Uber and Lyft, face the intricate real-time challenge of bundling and matching passengers—each with distinct origins and destinations—to available vehicles, all while navigating significant system uncertainties. Due to the extensive observation space arising from the large number of drivers and orders, order dispatching, though fundamentally a centralized task, is often addressed using Multi-Agent Reinforcement Learning (MARL). However, independent MARL methods fail to capture global information and exhibit poor cooperation among workers, while Centralized Training Decentralized Execution (CTDE) MARL methods suffer from the Curse of Dimensionality (CoD). To overcome these challenges, we propose Triple-BERT, a centralized framework designed specifically for large-scale order dispatching on ride-sharing platforms based on Single Agent Reinforcement Learning (SARL). Built on a variant TD3, our approach addresses the vast action space through an action decomposition strategy that breaks down the joint action probability into individual driver action probabilities. To handle the extensive observation space, we introduce a novel BERT-based network, where parameter reuse mitigates parameter growth as the number of drivers and orders increases, and the attention mechanism effectively captures the complex relationships among the large pool of driver and orders. We validate our method using a real-world ride-hailing dataset from Manhattan. Triple-BERT achieves approximately an 11.95% improvement over current state-of-the-art methods, with a 4.26% increase in served orders and a 22.25% reduction in pickup times. Our code, trained model parameters, and processed data are publicly available at the repository https://github.com/RS2002/Triple-BERT.  [^1] Do We Really Need MARL for Order Dispatch on Ride-Sharing Platforms? In summary: No, because we developed a SARL method that achieves better global planning. However, yes, our method still requires MARL for pre-training to establish a strong starting point for SARL.
@@ 1 Introduction
Recently, Reinforcement Learning (RL) methods have shown great potential in addressing the order dispatching problem in ride-sharing platforms. Model-free RL, in particular, enables agents to autonomously learn optimal dispatching policies by interacting with the environment, without requiring complex system modeling. This approach allows platforms to optimize multiple objectives, including platform income, driver payments, and customer satisfaction. Despite these advantages, applying RL to large-scale order dispatching introduces significant challenges. The vast action and observation spaces, stemming from the large number of drivers and orders, make sufficient exploration and efficient training difficult. Multi-Agent Reinforcement Learning (MARL) methods have been widely adopted to address these challenges by decomposing the problem into smaller subproblems for individual agents (drivers). Independent MARL methods, such as Independent Double DQN (IDDQN) [al2019deeppool; liu2022deep; wang2023optimization] and Independent SAC (ISAC) [zhang2024joint], are computationally efficient but fail to capture global information and exhibit limited cooperation among agents. Graph Neural Networks (GNNs) have been introduced to enable the network to capture neighboring information for each agent, alleviating this issue to certain extent [hu2025bmg; wang2024promoting]. Meanwhile, Centralized Training with Decentralized Execution (CTDE) methods, such as QMIX [hao2022hierarchical] and Coordinated Policy Optimization (CoPO) [wang2024dyps], struggle with the Curse of Dimensionality (CoD) when applied to large-scale scenarios with thousands of agents, resulting in slow convergence and suboptimal performance. (Due to page limitations, we provide a detailed review of ride-sharing methods and cooperative MARL approaches in Appendix F.)
@@
To address these limitations, this paper proposes a centralized Single-Agent Reinforcement Learning (SARL) method, named Triple-BERT, tailored for large-scale order dispatching in ride-sharing platforms. Triple-BERT introduces an action decomposition method that simplifies the joint action probability into individual driver action probabilities, enabling each driver to make independent decisions while maintaining global coordination. The method leverages TD3 [fujimoto2018addressing] for optimization, with modifications to the actor optimization process via policy gradient [sutton1999policy] to better suit the ride-sharing context. To handle the extensive observation space, we design a novel BERT-based [devlin2019bert] neural network architecture. This network employs bi-directional self-attention to effectively capture complex relationships between drivers and orders, while its parameter reuse mechanism prevents parameter explosion as the number of drivers and orders increases. Additionally, compared to MARL, SARL faces a unique challenge of sample scarcity, as the records of multiple agents are merged into a single training stream. To address this, we propose a two-stage training strategy, where feature extractors are pre-trained using a MARL approach to learn general embedding capabilities, followed by centralized fine-tuning. The main contributions of this paper can be summarized as follows:
@@
  • We introduce Triple-BERT, which is the first centralized framework for large-scale order dispatching on ride-sharing platforms based on a variant centralized TD3. This framework addresses the limitations of the observation space and the inefficiencies in cooperation among agents present in conventional MARL-based methods. To tackle the large action space inherent in the matching problem of order dispatching tasks, we propose an action decomposition method that breaks down the joint action probability into individual driver action probabilities. Additionally, we propose a two-stage training method to address the sample scarcity issue in SARL, where the feature extractors are first trained using a MARL approach.
@@ 2 Problem Setup
Figure 1: Workflow: At each time step, the worker and order pools update their states based on the assignments made in the previous time step. Specifically, the order pool adds newly arrived orders and removes overdue ones. For IDDQN, the Q-value of each worker-order pair is calculated, and ILP is applied to maximize the global Q-value. For TD3, the probability of each worker-order pair is computed, followed by the application of ILP to maximize the global assignment probability.
@@ 3 Methodology
Figure 2: Network Architecture: The network consists of three main components: the feature extractor, the actor sub-network, and the critic sub-network. First, a worker encoder and an order encoder are used to extract features from individual worker and order information, respectively. Then an Actor BERT model captures the relationships between them and a QK-Attention module calculates the selection probabilities for each worker-order pair. Finally, the fused features of the selected worker-order pairs are input into two separate Critic BERT models for further information extraction, and two Critic MLPs compute the Q-values, as TD3 requires two critics. (In this figure, the fused sequence (input to Critic-BERT) represents workers $1$, 33, 66, and $n$ selecting orders $2$, 33, 44, and $m$, respectively.)
@@ 3.1 Overview
In this work, we aim to utilize centralized SARL to address the large-scale order dispatching task, with the goal of enabling the model to fully leverage global information to enhance cooperation among workers. To tackle the challenges of large action and observation spaces, we propose a novel network architecture, as illustrated in Fig. 2. This architecture employs the BERT model [devlin2019bert] to effectively extract the relationships between workers and orders using the self-attention mechanism. Additionally, an improved QK-attention [zhao2025crossfi] is implemented to reduce the computational complexity associated with the order dispatching task. Furthermore, we introduce an action decomposition method that breaks down the choice probability of each action within the vast action space into individual action probabilities for each worker selecting each order. Finally, to address the data scarcity challenge in MARL, we propose a two-stage training method, as shown in Fig. 1. In the first stage, we train the upstream layers of the network using the IDDQN approach, allowing them to develop general feature extraction capabilities. Subsequently, we train the entire neural network using centralized TD3 to realize better cooperation between workers.
@@ 3.2.1 Feature Extractors
At each time step, the network takes the entire state $S_{t}=[W_{t},O_{t}]$ as input. We consider this as a combination of two sequences: $W_{t}$ and $O_{t}$. For each element $w_{i,t}$ and $o_{j,t}$, we employ two distinct encoders, referred to as the “Worker Encoder” and the “Order Encoder”, to embed them separately into a feature space of the same dimension, allowing them to be input into a single BERT model.
Each worker state $w_{i,t}$ consists of two parts: an on-board order sequence and other non-sequence information. For the order sequence, a bi-directional LSTM [LSTM] is utilized to extract its features. This approach effectively encodes variable-length sequences into a uniform dimensional feature space, addressing the CoD associated with conventional MLP encoders, where the number of parameters increases with sequence length. For non-sequence information, an MLP is employed for feature extraction. Finally, the two features are combined into a primary feature $\tilde{w}_{i,t}$. For the orders to be assigned $o_{j,t}$, an MLP is also used to extract the feature $\tilde{o}_{j,t}$. Notably, the dimensions of $\tilde{w}_{i,t}$ and $\tilde{o}_{j,t}$ are identical, and their information is concatenated into a sequence represented as $\tilde{S}_{t}=[\tilde{w}_{1,t},\tilde{w}_{2,t},\ldots,\tilde{w}_{n,t},\tilde{o}_{1,t},\tilde{o}_{2,t},\ldots,\tilde{o}_{m_{t},t}]$.
Additionally, to facilitate network convergence and enhance the extraction of input features, we incorporate a normalization layer and an Adaptive Re-weighting Layer (ARL) [chen2024modelling]. Given that different parts of the input may have varying magnitudes, which can impede model training, the normalization layer effectively addresses this issue. Furthermore, since different parts of the input carry different levels of importance, we utilize the ARL to enable the model to learn these variations, represented as: $y=x\circ\Omega$, where $x$ denotes the input, $\Omega$ represents the weight vector, calculated by $\Omega=\text{MLP}(x)$, and $\circ$ indicates the element-wise product.
@@ 3.2.2 Actor Sub-Networks
The Actor sub-network consists of a BERT [devlin2019bert] model for feature extraction and a QK-attention module [zhao2025crossfi] for action decomposition and generation, which we will introduce in turn. In the feature extractors, we have already extracted the primary features from each worker and order state separately. To further explore the relationships between workers and orders, we utilize the BERT model, where the self-attention mechanism can effectively capture these relationships: $\overline{S}_{t}=[\overline{w}_{1,t},\overline{w}_{2,t},\ldots,\overline{w}_{n,t},\overline{o}_{1,t},\overline{o}_{2,t},\ldots,\overline{o}_{m_{t},t}]=\text{Actor-BERT}(\tilde{S}_{t})$. Specifically, due to the permutation invariance of our input sequence, we omit the positional embedding in BERT, ensuring that the order in $S$ does not influence the encoding result. In contrast to conventional MARL methods like [hu2025bmg; hao2022hierarchical], which encode each worker with its neighboring states to gain a broader perspective, our Actor-BERT directly aggregates global worker information, facilitating more effective cooperative dispatching between workers.
In conventional order dispatching tasks, the typical approach to address the dynamic action space (related to the number of orders) involves evaluating each worker-order pair separately and finding the optimal dispatching solution based on these evaluations. However, this approach has two significant shortcomings. First, it neglects the relationships between orders, which we address through the self-attention mechanism in BERT, capturing not only the relationships between workers but also between orders and between orders and workers. Second, evaluating each worker-order pair is time-consuming and resource-intensive: $\text{F}(\overline{w}_{i,t},\overline{o}_{j,t};\theta_{F})\in\mathbb{R}^{1}$, where F is the network and $\theta_{F}$ represents its parameters. The complexity can be represented as $O(|\text{F}|\cdot n\cdot m_{t})$, where $|\text{F}|$ denotes the complexity of the neural network. To mitigate this issue, we employ a QK-attention module [zhao2025crossfi], represented as:
@@
where f and g are two smaller networks, and $\theta_{f}$ and $\theta_{g}$ are their parameters. The intuition behind QK-attention is to use two smaller networks to approximate a larger network, similar to the motivation behind LoRA [hu2022lora]. In this way, the complexity of computing all worker-order pairs becomes $O(|\text{f}|\cdot n+|\text{g}|\cdot m_{t}+d\cdot n\cdot m_{t})$, where $|\text{f}|$ and $|\text{g}|$ are the complexities of the two neural networks, $d$ is their output dimension, and $d\cdot n\cdot m_{t}$ is the complexity of matrix multiplication. Here, $d$ is very small, making $d\cdot n\cdot m_{t}$ much smaller than the neural network computation complexity, i.e., $d\cdot n\cdot m_{t}\ll|\text{f}|\approx|\text{g}|<|\text{F}|$. Thus, we have $O(|\text{f}|\cdot n+|\text{g}|\cdot m_{t}+d\cdot n\cdot m_{t})<O(|\text{F}|\cdot(n+m_{t}))<O(|\text{F}|\cdot n\cdot m_{t})$, indicating that the QK-attention successfully transforms the multiplication complexity of evaluating each worker-order pair into addition complexity.
However, we observe a parameter redundancy issue in Equation 2, which can lead to potential instability during training. This redundancy arises because there are actually infinite solutions for $f$ and $g$, as $f^{\prime}=\alpha f$ and $g^{\prime}=\frac{g}{\alpha}$ is also a valid solution for any non-zero real vector $\alpha$. Inspired by Dueling DQN [wang2016dueling], we propose a positive normalization method:
@@ 3.2.3 Critic Sub-Networks
where $\overline{w}_{i,t}$ and $\overline{o}_{j,t}$ are the outputs of Actor-BERT, and $(\overline{w}_{i,t},\overline{o}_{j,t})$ represents the combination of the two vectors into a single feature vector. We then construct a new sequence: $\dot{S}_{t}=[\mathcal{A}(w_{1,t}),\mathcal{A}(w_{2,t}),\ldots,\mathcal{A}(w_{i,t})]$. Another BERT network, referred to as “Critic-BERT”, is used to further extract features from $\dot{S}_{t}$, represented as $\ddot{S}_{t}=\text{Critic-BERT}(\dot{S}_{t})$. A self-attention mechanism and a linear layer (collectively named Critic-MLP) are then utilized to estimate the Q-value from $\ddot{S}_{t}$ (for detailed processing methods, refer to [chou2024midibert]). Furthermore, as TD3 [fujimoto2018addressing] requires two critics, we employ two distinct Critic-BERT and Critic-MLP networks. These share the input features from Actor-BERT but process them separately.
@@ 3.3.1 Stage 1: Decentralized IDDQN Training
To construct the network, we utilize QK-attention to process the outputs of the worker encoder and order encoders to estimate the Q-value for each worker-order pair, represented as $\text{QK-Attention-Norm}(\tilde{w}_{i,t},\tilde{o}_{j,t})$ (denoted as $y_{i,j,t}$). Although the state space encompasses the entire order state from $o_{1,t}$ to $o_{m_{t},t}$, we focus on a single order $o_{j,t}$ when computing the Q-value for choosing order $j$. This approach aligns with previous work such as [hu2025bmg; hu2025coordinating], as the entire order state can be excessively large for a simple network to learn (our Triple-BERT effectively addresses this issue) and many networks struggle to process variable dimensional inputs (with order amounts varying at each time step). Consequently, we can compute a Q-matrix $Y_{t}\in\mathbb{R}^{n,m_{t}}$, where the element in the $i$-th row and $j$-th column, $y_{i,j,t}$, represents the Q-value of assigning order $j$ to worker $i$ at time $t$. The core strategy of IDDQN is to maximize the global Q-value, expressed as $\text{Q}(S_{t},A_{t})=\sum_{i=1}^{n}\text{Q}(s_{i,t},a_{i,t})$ at each time step. To achieve this, we construct a bipartite graph where each worker and order is represented as a node. An arbitrary worker $i$ and order $j$ are linked by an edge weighted by the Q-value of this worker selecting this order at the current time, i.e., $y_{i,j,t}$. We then utilize Integer Linear Programming (ILP) to solve this maximizing bipartite matching problem. (To avoid assigning orders to unavailable workers—those at full capacity or on their way to pick up an assigned order—we set the Q-value of all actions for such workers in the Q-matrix $Y_{t}$ to -$\infty$.) A detailed construction of the problem is provided in Appendix B.1. For the training of IDDQN, it follows the same process of previous work [hu2025bmg]. Due to page limitation, we detailed it in Appendix D.1.
Here, We want to make some explanations about the independent assumption. We acknowledge that the independent assumption can be unreliable, which indeed hinders the performance of conventional independent MARL-based methods. However, previous works [shi2019operating; feng2022coordinating; wang2023reassignment] have shown the efficacy and simplicity of independent MARL methods. As a result, even if the performance of independent MARL is not satisfactory, it does provide a good starting point for our centralized SARL method. In our approach, the independent assumption is only utilized during the pre-training stage to warm up the model. After pre-training, our centralized SARL framework no longer relies on the independence assumption.